OpenStudy (anonymous):

3x-y=-6 what is the slope intercept form of this?

OpenStudy (anonymous):

Slope-intercept form is in the form of y = mx + c. So just rearrange and put the y on the left and everything else on the right.

OpenStudy (anonymous):

y=3x-6?

OpenStudy (anonymous):

It's + 6 not - 6

OpenStudy (anonymous):

thankyou and i have more can you help

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

x+3y=-6

OpenStudy (anonymous):

It's the same process

OpenStudy (anonymous):

3y=x+6?

OpenStudy (anonymous):

No, it seems you are confused about rearranging the terms

OpenStudy (anonymous):

a + b = c a + b - a = c - a b = c - a

OpenStudy (anonymous):

so would it be 3y=-6+x

OpenStudy (anonymous):

The x needs to be negative as well x + 3y = -6 x + 3y - x = -6 - x 3y = -6 - x

OpenStudy (anonymous):

But the form is in y = mx + c and you got 3y. So you need to divide the equation by 3